OpenStudy (anonymous):

Part A: Solve -np-80 < 60 for n. Show your work Part B: Solve 2a-5d=30 for d show your work

OpenStudy (anonymous):

@pooja195 help me please

OpenStudy (jjwalton04):

A: n>-140/p B: d=2/5a-6

OpenStudy (jjwalton04):

hope im correct

OpenStudy (anonymous):

@jjwalton04 can you show me your work?

OpenStudy (jjwalton04):

Let's solve for d. 2a−5d=30 Step 1: Add -2a to both sides. 2a−5d+−2a=30+−2a −5d=−2a+30 Step 2: Divide both sides by -5. -5d/5=-2a+30/5 ---------------------------------------------- Let's solve for n. (−n)(p)−80<60 Step 1: Add 80 to both sides. −np−80+80<60+80 −np<140 Step 2: Divide both sides by -p -np/-p <140/-p

OpenStudy (jjwalton04):

so am i right